OpenStudy (thomasfire55):

Choose the function that has the same graph as y = -4 cos(x - π).

OpenStudy (thomasfire55):

y = -4sin(x - π) y = -4sin(x - (π/2)) y = -4sin(x + π) y = -4sin(x + (π/2))

OpenStudy (thomasfire55):

@agent0smith

OpenStudy (agent0smith):

Easy way to do it is to graph them...

OpenStudy (thomasfire55):

@Hannahsayshi

OpenStudy (hannahsayshi):

I'm giving that guy the best response because he's exactly right. If you graph it you should come up with the answer. It won't be that hard.

OpenStudy (thomasfire55):

ugh so much work

OpenStudy (hannahsayshi):

I'm sorry I was told I cannot give direct responses.

OpenStudy (thomasfire55):

wow

OpenStudy (hannahsayshi):

All you have to do is graph it correctly and it'll come up.

OpenStudy (thomasfire55):

desmos to the rescue